OpenStudy (anonymous):

need help please if i dont get this right i wont pass!!!! What is the second term of this multiplication pattern? multiplication pattern using the formula 5 • 5n – 1

OpenStudy (mathmale):

I'm pretty sure you mean this:\[5 • 5^{n-1} \]

OpenStudy (anonymous):

yeah

OpenStudy (mathmale):

Does this look like the original problem? 1) Let n-1. Evaluate \[5 • 5^{n-1} ~for~n=1\]

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

What is the value?

OpenStudy (mathmale):

\[5 • 5^{1-1}=?\]

OpenStudy (anonymous):

i dont know what it equals

OpenStudy (mathmale):

Hint:\[5^{1-1}\] is not zero.

OpenStudy (mathmale):

What is 5^0?

OpenStudy (anonymous):

is it just 25?

OpenStudy (mathmale):

\[5^0=?\]

OpenStudy (anonymous):

0

OpenStudy (mathmale):

No. \[5^2=25.~ 5^1=?\]

OpenStudy (mathmale):

"Five to the first power is ... ??"

OpenStudy (anonymous):

1 sec gotta turn my lamp on

OpenStudy (anonymous):

ok i think it is 25?

OpenStudy (mathmale):

If you have a TI calculator, type in 5^1 ENTER

OpenStudy (anonymous):

5

OpenStudy (mathmale):

If you have some other calculator, try typing in this:\[5 ~y^x 1~ ENTER\]

OpenStudy (anonymous):

i have a ti i just had to find it. it says it's 5